# Preparation of high fidelity entangled cat states with composite pulses

### v1

@@ Abstract
We propose a protocol for the preparation of high-fidelity entangled cat states with composite pulses. The physical model contains two Kerr-nonlinear resonators and a cavity. By properly designing the parameters, each Kerr-nonlinear resonator is confined in the cat-state subspace and the entangled cat states can be generated efficiently. We introduce composite two-photon drives with multiple amplitudes and frequencies to improve the fidelity of the entangled cat states in the presence of parameter errors. The performance of the protocol is estimated by taking into account the parametric errors and decoherence. Numerical simulation results show that, the protocol is insensitive to timing error and detuning error, and has strong robustness to decoherence. We hope the protocol may provide a method for preparing stable entangled cat states.
@@ I Introduction
Cat-state qubits are one of the promising bonsic quantum qubits in quantum information processing Ralph et al. (2003); Zhou et al. (2021); Shen et al. (2015); Zheng (1998); Li et al. (2017); Chen et al. , whose two logical qubits are usually represented by two orthogonal cat states (i.e., superposition of coherent states) Gilchrist et al. (2004); Chen et al. (2021); Kang et al. (2022); Liu et al. (2022). Cat-state qubits have many unique advantages. For instance, they are noise biased Ma et al. (2021); Cai et al. (2021). The phase-flip error can be effectively suppressed, only the bit-flip error needs to be concerned in error correction. Thus, the number of building blocks for error correction can be significantly reduced Guillaud and Mirrahimi (2019); Mirrahimi et al. (2014); Puri et al. (2019); Chen et al. (2022). In addition, cat qubits have an enhanced lifetime with error corrections Ofek et al. (2016); Li et al. (2024). Therefore, cat-state qubits have received much attention and many protocols in view of cat-state qubits have been proposed, such as nonadiabatic geometric quantum computation with cat-state qubits Kang et al. (2022) and one-step parity measurement of $N$ cat-state qubits Li et al. (2024). Entangled cat states are also important for the demonstration of the fundamentals of quantum physics, and have wide applications in modern quantum technologies Chen et al. (2022, 2021). Protocols have been presented for producing entangled cat states, such as the generation of giant entangled cat states Chen et al. (2021) and the preparation of multidimensional entangled cat states in cavity quantum electrodynamics (QED) Yang et al. (2023); Zhang et al. (2020).
@@
In this paper, we consider to use the ${\o}$${\o}$rensen (MS) entangling gate for the preparation of entangled cat states Sørensen and Mølmer (1999, 2000). The MS entangling gate has some specific characteristics. Firstly, it makes qubits independent of the decoherence of the motion mode Sørensen and Mølmer (1999, 2000); Mitra et al. (2020). Secondly, it possesses a built-in noise-resilience feature against certain types of local noise Chen et al. (2022). Therefore, the MS entangling gate has become an important resource and inspired a wide range of research interests Zhu et al. (2006); Choi et al. (2014); Green and Biercuk (2015); Ballance et al. (2016); Gaebler et al. (2016); Manovitz et al. (2017); Mitra et al. (2020); Brickman et al. (2005). For example, in 2021, Chen $et$ $al$. proposed that entangled cat states can be prepared by using cat-code MS gate Chen et al. (2022). However, the protocol Chen et al. (2022) could be sensitive to timing error. We note that some kind of composite pulses can play a role in robust control of quantum system and reduce errors Haddadfarshi and Mintert (2016); Webb et al. (2018); Shapira et al. (2018); Wang et al. (2021); Palmero et al. (2017); Manovitz et al. (2017); Leung et al. (2018). For example, Haddadfarshi $et$ $al$. pointed out that properly designed polychromatic control pulses can greatly reduce errors Haddadfarshi and Mintert (2016) and then Webb $et$ $al$. demonstrated the protocol experimentally Webb et al. (2018). Shapira $et$ $al$. generalized the MS entangling gate by using additional frequency components in the laser drive Shapira et al. (2018). Wang $et$ $al$. used suitably designed laser pulse with modulated amplitude and phase to improve the fidelity of the MS entangling gate Wang et al. (2021).
@@
Inspired by the literatures above in this paper, we propose to use cat-code MS gate to prepare high-fidelity entangled cat states. We introduce a time-independent drive which is constructed by composite two-photon drives with different amplitudes and frequencies. Under such a drive, we can obtain more stable entangled cat states with appropriate parameters. Meanwhile, the sensitivities of the entangled cat states against errors are significantly reduced, and the robustness to decoherence are obviously improved.
@@ II Cat-code MS entangling gate
The physical model consists of two Kerr-nonlinear cavities ($\mathcal{A}_{1},\mathcal{A}_{2}$) with the same frequency $\omega$, and another cavity ($\mathcal{A}_{0}$) with frequency $\omega_{0}$. The Hamiltonian of this system in the interaction picture is given by (assume $\hbar=1$)
@@
where $J$ is the coupling strength, $a_{k}(k=1,2)$ is the annihilation operator of the Kerr-nonlinear resonator, $a_{0}$ is the annihilation operator of the cavity, and $\Delta=\omega_{0}-\omega$ is the detuning. Each Kerr-nonlinear resonator is resonantly driven by a single-mode two-photon squeezing drive with frequency $\omega_{p}=2\omega$ and amplitude $\Omega_{p}$. The Hamiltonian of the Kerr parametric oscillators (KPOs) $H_{k}^{\mathrm{Kerr}}$ in Eq. (1) is Puri et al. (2019, 2017)
@@
where $K$ is Kerr nonlinearity coefficient Miranowicz et al. (2016); Bourassa et al. (2012). To construct cat-state qubits, the Hamiltonian $H_{k}^{\mathrm{Kerr}}$ in Eq. (3) can be rewritten as:
@@
Equation (4) clearly shows that when $\alpha=\sqrt{\frac{\Omega_{p}}{K}}$, the two coherent states |±$|\pm\alpha\rangle$ are degenerate eigenstates of the annihilation operator $a_{k}$. Therefore, the superpositions of the coherent states Cai et al. (2021); Puri et al. (2019, 2017, 2020)
@@
The first term in the right-hand side of Eq. (9) can be dropped because it is the identity matrix $(|C_{-}\rangle_{k}\langle{C_{-}}|+|C_{+}\rangle_{k}\langle{C}_{+}|)$ of the subspace. We define Pauli matrices $\sigma^{+}_{k}=|C_{-}\rangle_{k}\langle C_{+}|$ and $\sigma^{-}_{k}=|C_{+}\rangle_{k}\langle C_{-}|$. Then the Hamiltonian $H_{\mathrm{eff}}$ in Eq. (9) reduces to
@@
To well understand the evolution of the system, the exact solution of the evolution operator based on the effective Hamiltonian $H_{\mathrm{eff}}$ can be calculated with the Magnus expansion Blanes et al. (2010, 2009):
@@
Note that a large detuning $\Delta$ is necessary for the Magnus expansion to proceed. Setting $t=T=2\pi/\Delta$, there has $\chi(T)=0$ and $\beta(T)=(\frac{2J\alpha}{\Delta})^{2}\cdot(-2\pi)$. When choosing the parameter $\Delta=4J\alpha$, then $\beta(T)=-\frac{\pi}{2}$. Thus the evolution operator at time $T$ reads
@@
The $U(T)$ in Eq. (15) is the well-known MS entangling gate Sørensen and Mølmer (2000). The MS entangling gate is an important physical resource in quantum computing, for example, it can be applied to the steps of marking and state amplification in Grover’s quantum search algorithm Brickman et al. (2005); Ohno (2024); Yang et al. (2021); Sakhouf et al. (2021); Anikeeva et al. (2021); Lu et al. (2022). Under the action of $U(T)$, we can realize the preparation of the entangled cat states. As shown in Eq. (16), the left side of Eq. (16) are the four different initial states (the input states), the right side are the corresponding maximum entangled cat states (the output state) obtained via $U(T)$ Chen et al. (2022); Sørensen and Mølmer (1999).
@@
For convenience, the initial state of the system is set as $|\Psi_{in}\rangle=|C_{+}\rangle|C_{+}\rangle|0\rangle$, and the corresponding entangled cat state is $|\Psi_{out}\rangle=\frac{1}{\sqrt{2}}(|C_{+}\rangle|C_{+}\rangle|0\rangle+i|C_{-}\rangle|C_{-}\rangle|0\rangle)$ in this manuscript. All numerical simulation results in this section are obtained by using the full Hamiltonian given by Eq. (1). Here, the fidelity of entangled state is given by $F=\langle\Psi|\rho|\Psi\rangle$, where $|\Psi\rangle$ is the target entangled cat state i.e., $|\Psi_{out}\rangle$, and $\rho$ is the density operator. The experimental parameters are chosen as $\alpha=2$, the Kerr nolinearity $K/2\pi=20\mathrm{MHz}$ Wang et al. (2019), $\Omega_{p}=K\alpha^{2}$, $J=2\pi\mathrm{MHz}$, and $\Delta=4J\alpha$.
Figure 2: (a) Population envolution of the cat states. The solid line represents the state $|C_{+}\rangle|C_{+}\rangle|0\rangle$, and the dotted line represents the state $|C_{-}\rangle|C_{-}\rangle|0\rangle$. (b) The fidelity against timing error $\delta_{t}$, error range is selected for [-0.1, 0.1]. The experimental parameters are set as $\alpha=2$, $T=2\pi/\Delta$, the Kerr nolinearity $K/2\pi=20\mathrm{MHz}$, the driving amplitude $\Omega_{p}=K\alpha^{2}$, $J=2\pi\mathrm{MHz}$, and $\Delta=4J\alpha$.
In Fig. 2, we plot the time evolutions of the states $|C_{+}\rangle|C_{+}\rangle|0\rangle$ and $|C_{-}\rangle|C_{-}\rangle|0\rangle$. As is shown, the populations of both states are close to $50\%$ around the gate time $T$. This means that the protocol is feasible, and the maximum entangled cat state can be successfully prepared when $t=T$. However, the high-fidelity entangled cat state can only exist when the time is extremely close to $T$. When the timing error $\delta_{t}$ deviates from 0, the fidelity of the entangled cat state $|\Psi_{out}\rangle$ drops significantly, which can be seen in Fig. 2. Here, the relative error of the parameter $x$ is defined via $\delta_{x}=(x^{\prime}-x)/x$, in which $x^{\prime}$ denotes the actual value and $x$ is the ideal value.
@@ III Preparation of high-fidelity entangled cat state
The last term in Eq. (24) is the additional composite drives, and $\delta_{n}=\omega_{n}-\omega$ is the detuning of the $n$th two-photon drive. According to effective Hamiltonian theory James and Jerke (2007), the effective Hamiltonian in the limit of large detunings can be written as
@@
where $\frac{1}{\varepsilon_{n}}=({\frac{1}{\Delta}+\frac{1}{\delta_{n}}})$. Similar to the derivation from Eq. (1) to Eq. (10), we can get a new effective Hamiltonian by restricting the Hamiltonian $H_{\mathrm{eff},N}(t)$ to the cat-state subspace
@@
For simplicity, we define $\frac{4\Omega_{n}}{\varepsilon_{n}}=r_{n}$ and $\zeta=(\Delta-\delta_{n})/n$ with $n=1,2,3,...,N$. Then the effective Hamiltonian $H_{\mathrm{eff},N}^{\prime}(t)$ can be expressed as
@@
To understand the construction of composite pulses, we introduce the dimensionless position operator $x=\frac{1}{\sqrt{2}}(a_{0}+a_{0}^{{\dagger}})$ and the momentum operator $p=\frac{i}{\sqrt{2}}(a_{0}^{{\dagger}}-a_{0})$. The effective Hamiltonian in Eq. (29) can be reformulated as
$\mathcal{H}_{\mathrm{MS},N}(t)=f(t)S_{x}x+g(t)S_{x}p.$ (30)
@@
As seen in Eq. (31), the evolution of the operator $\mathcal{U}(t)$ follows the trajectory with dimensionless coordinates $\mathcal{F}(t)$ and $\mathcal{G}(t)$ in the $x-p$ phase-space. Equation (32) shows that $\mathcal{F}(t)$ and $\mathcal{G}(t)$ are respectively the integrations of $f(t)$ and $g(t)$ Sørensen and Mølmer (2000); Haddadfarshi and Mintert (2016). Ideally, when the time $t=\tau=\frac{2\pi}{\zeta}$, $\mathcal{F}(\tau)=0$, $\mathcal{G}(\tau)=0$, and the phase-space trajectory of $(\mathcal{F}(t),\mathcal{G}(t))$ is closed as seen in Fig. 3(a). Meanwhile, the operator $\mathcal{U}(\tau)=e^{-i\mathcal{A}(\tau)S_{x}^{2}}$, that is, only the internal spin evolution is retained. However, errors will lead to the phase-space trajectory not closed completely, which in turn can causes infidelity of entangled cat states Rauschenbeutel et al. (1999); Wang et al. (2021); Webb et al. (2018). We can see from Fig. 3(b), the circular trajectory becomes more complete with the increase of $N$. In other words, we can reduce errors of the entangled cat states by increasing $N$.
@@
In Fig. 4, we plot the population evolutions of the cat states with different $N$ ($N=1,2,4,8$). We can see from Fig. 4 that the curves become more flat near time $T$ with the increase of $N$. This means that more stable entangled cat state can be obtained by simply increasing $N$. Here, the parameters $\{\alpha,K,\Omega_{p},\Delta\}$ are same as Fig. 2, and $\zeta=\Delta$.
Figure 5: Fidelity versus parameter errors. (a) timing error, (b) detuning error, (c) coupling error. In figures (a) and (b), increasing $N$ leads to a flatter response to control errors, the robustnesses against parameters errors are enhanced. The other parameters are the same as in Fig. 4

### v2

# Preparation of high-fidelity entangled cat states with composite pulses

@@ Abstract
We propose a protocol for the preparation of high-fidelity entangled cat states with composite pulses. The physical model contains two Kerr-nonlinear resonators and a cavity. By properly designing the parameters, each Kerr-nonlinear resonator is confined in the cat-state subspace and the entangled cat states can be generated efficiently. We introduce composite two-photon drives with multiple amplitudes and frequencies to improve the fidelity of the entangled cat states in the presence of parameter errors. The performance of the protocol is estimated by taking into account the parametric errors and decoherence. Numerical simulation results show that, the protocol is robustness to timing error, detuning error, and decoherence. We hope the protocol may provide a method for preparing stable entangled cat states.
@@ I Introduction
Cat-state qubits are one kind of the promising bonsic quantum qubits in quantum information processing Ralph et al. (2003); Zhou et al. (2021); Shen et al. (2015); Zheng (1998); Li et al. (2017), whose two logical qubits are usually represented by two orthogonal cat states (i.e., superposition of coherent states) Gilchrist et al. (2004); Chen et al. (2021); Kang et al. (2022); Liu et al. (2022). Cat-state qubits have some unique advantages. For instance, they are noise biased Ma et al. (2021); Cai et al. (2021). The phase-flip error can be effectively suppressed, only the bit-flip error needs to be concerned in error correction. Thus, the number of building blocks for error correction can be significantly reduced Guillaud and Mirrahimi (2019); Mirrahimi et al. (2014); Puri et al. (2019); Chen et al. (2022). In addition, cat qubits have an enhanced lifetime with error corrections Ofek et al. (2016); Li et al. (2024). Therefore, cat-state qubits have received much attention and many protocols Kang et al. (2022); Li et al. (2024); Kang et al. (2022); Chen et al. (2024); Etesse et al. (2015); Zhang et al. (2021); Wang et al. ; Ourjoumtsev ; Han et al. in view of cat-state qubits have been proposed, such as experimental preparation squeezed cat states Wang et al. , one-step parity measurement of $N$ cat-state qubits Li et al. (2024) and the remote preparation of cat state Han et al. . Entangled cat states are also important for the demonstration of the fundamentals of quantum physics, and have wide applications in modern quantum technologies Chen et al. (2022); Yang et al. (2023); Zhang et al. (2020); Chen et al. (2021, 2021); Yang and Zheng (2018). Protocols have been presented for producing entangled cat states, such as the generation of giant entangled cat states Chen et al. (2021) and the preparation of multidimensional entangled cat states in cavity quantum electrodynamics (QED) Yang et al. (2023); Zhang et al. (2020).
@@
Among the various methods for preparing entangled cat states, the ${\o}$${\o}$rensen (MS) entangling gate is a commonly employed technique Chen et al. (2021, 2022). The MS entangling gate has some specific characteristics Sørensen and Mølmer (1999, 2000). Firstly, it makes qubits independent of the decoherence of the motion mode Sørensen and Mølmer (1999, 2000); Mitra et al. (2020). Secondly, it possesses a built-in noise-resilience feature against certain types of local noise Chen et al. (2022). Therefore, the MS entangling gate inspired a wide range of research interests Zhu et al. (2006); Choi et al. (2014); Green and Biercuk (2015); Ballance et al. (2016); Gaebler et al. (2016); Manovitz et al. (2017); Mitra et al. (2020); Brickman et al. (2005). For example, in 2021, Chen $et$ $al$. used cat-code MS gate to prepare entangled cat states Chen et al. (2022). But the protocol Chen et al. (2022) could be sensitive to timing error. We note that some kinds of composite pulses can play a role in robust control of quantum system and reduce errors Haddadfarshi and Mintert (2016); Webb et al. (2018); Shapira et al. (2018); Wang et al. (2021); Palmero et al. (2017); Manovitz et al. (2017); Leung et al. (2018). For example, Haddadfarshi $et$ $al$. pointed out that properly designed polychromatic control pulses can greatly reduce errors Haddadfarshi and Mintert (2016) and then Webb $et$ $al$. demonstrated the technique experimentally Webb et al. (2018). Shapira $et$ $al$. generalized the MS entangling gate by using additional frequency components in the laser drive Shapira et al. (2018). Wang $et$ $al$. used suitably designed laser pulse with modulated amplitude and phase to improve the fidelity of the MS entangling gate Wang et al. (2021).
@@
Inspired by the literatures above in this paper, we propose to use cat-code MS gate and composite pulses to prepare high-fidelity entangled cat states. The composite pulses are constructed by two-photon drives with different amplitudes and frequencies. Under such a drive, we can obtain more stable entangled cat states with appropriate parameters. Meanwhile, the sensitivities of the entangled cat states against errors are significantly reduced, and the robustness to decoherence are obviously improved.
@@ II Cat-code MS entangling gate
The physical model consists of two Kerr-nonlinear cavities ($\mathcal{A}_{1},\mathcal{A}_{2}$) with the same frequency $\omega_{k}$, and another cavity ($\mathcal{A}_{0}$) with frequency $\omega_{0}$. The Hamiltonian of this system in the interaction picture is given by (assume $\hbar=1$)
@@
where $J$ is the coupling strength, $a_{k}(k=1,2)$ is the annihilation operator of the Kerr-nonlinear resonator, $a_{0}$ is the annihilation operator of the cavity, and $\Delta=\omega_{0}-\omega_{k}$ is the detuning. Each Kerr-nonlinear resonator is resonantly driven by a single-mode two-photon squeezing drive with frequency $\omega_{p}=2\omega_{k}$ and amplitude $\Omega_{p}$. The Hamiltonian of the Kerr parametric oscillators (KPOs) $H_{k}^{\mathrm{Kerr}}$ in Eq. (1) is Puri et al. (2019, 2017)
@@
where $K$ is Kerr nonlinearity coefficient Miranowicz et al. (2016); Bourassa et al. (2012). The Hamiltonian $H_{k}^{\mathrm{Kerr}}$ can be rewritten as:
@@
Equation (4) clearly shows that when $\alpha=\sqrt{\frac{\Omega_{p}}{K}}$, the two coherent states |±$|\pm\alpha\rangle$ are degenerate eigenstates of the annihilation operator $a_{k}$. Therefore, the superpositions of the coherent states Cai et al. (2021); Puri et al. (2019, 2017, 2020); Frattini et al. (2024)
@@
The first term in the right-hand side of Eq. (9) can be dropped because it is the identity matrix of the cat-state subspace. Defining Pauli matrices $\sigma^{+}_{k}=|C_{-}\rangle_{k}\langle C_{+}|$ and $\sigma^{-}_{k}=|C_{+}\rangle_{k}\langle C_{-}|$, the Hamiltonian in Eq. (9) reduces to Chen et al. (2022); Grimm et al. (2020); Ding et al. (2024)
@@
To well understand the evolution of the system, the evolution operator based on the effective Hamiltonian $H_{\mathrm{eff}}$ can be calculated with the Magnus expansion (see Appendix A for details) Magnus (1954); Blanes et al. (2010, 2009):
@@
Note that a large detuning $\Delta$ is necessary for the Magnus expansion to proceed. Setting $t=T=2\pi/\Delta$, there have $\chi(T)=0$ and $\beta(T)=(\frac{2J\alpha}{\Delta})^{2}\cdot(-2\pi)$. When choosing the parameter $\Delta=4J\alpha$, then $\beta(T)=-\frac{\pi}{2}$. Thus the evolution operator at time $T$ reads
@@
The $U(T)$ in Eq. (15) is the well-known MS entangling gate Sørensen and Mølmer (2000). The MS entangling gate is an important physical resource in quantum computing, for example, it can be applied to prepare entangled states Chen et al. (2022); Sørensen and Mølmer (1999). Under the action of $U(T)$, we can realize the preparation of the entangled cat states. As shown in Eq. (16), the left side of Eq. (16) are the four different initial states (the input states), and the right side are the corresponding maximum entangled cat states (the output states) obtained via $U(T)$ Chen et al. (2022); Sørensen and Mølmer (1999).
@@
For convenience, the initial state of the system is set as $|\Psi_{in}\rangle=|C_{+}\rangle|C_{+}\rangle|0\rangle$, and the corresponding entangled cat state is $|\Psi_{out}\rangle=\frac{1}{\sqrt{2}}(|C_{+}\rangle|C_{+}\rangle|0\rangle+i|C_{-}\rangle|C_{-}\rangle|0\rangle)$ in this manuscript. All numerical simulation results in this section are obtained by using the full Hamiltonian given by Eq. (1). Here, the fidelity of entangled state is given by $F=\langle\Psi|\rho|\Psi\rangle$, where $|\Psi\rangle$ is the target entangled cat state i.e., $|\Psi_{out}\rangle$, and $\rho$ is the density operator. The experimental parameters are chosen as $\alpha=2$, $K/2\pi=20\mathrm{MHz}$, $\Omega_{p}=K\alpha^{2}$, $J=2\pi\mathrm{MHz}$, and $\Delta=4J\alpha$ Chen et al. (2022).
Figure 2: (a) Population envolution of the cat states. The solid line represents the state $|C_{+}\rangle|C_{+}\rangle|0\rangle$, and the dotted line represents the state $|C_{-}\rangle|C_{-}\rangle|0\rangle$. (b) The fidelity against timing error $\delta_{t}$, error range is selected for [-0.1, 0.1]. The experimental parameters are set as $\alpha=2$, $T=2\pi/\Delta$, the Kerr nolinearity $K/2\pi=20\mathrm{MHz}$, the coupling strength $J/2\pi=1\mathrm{MHz}$ the driving amplitude $\Omega_{p}=K\alpha^{2}$, and the detuning $\Delta$ should obey $\Delta=4J\alpha$ Chen et al. (2022).
In Fig. 2, we plot the population evolutions with time of the states $|C_{+}\rangle|C_{+}\rangle|0\rangle$ and $|C_{-}\rangle|C_{-}\rangle|0\rangle$. As it is shown, the populations of both states are close to $50\%$ around the gate time $T$. That is, the maximum entangled cat state can be successfully prepared when $t=T$. However, the maximum entangled cat state can only exist when time is extremely close to $T$. To clearly see the fidelity of the maximum entangled cat state against timing errors, we plot Fig. 2. When the timing error $\delta_{t}$ deviates from 0, i.e., the actual operation time deviates from $T$, the fidelity of the entangled cat state $|\Psi_{out}\rangle$ drops significantly. Here, the relative error of the parameter $x$ is defined via $\delta_{x}=(x^{\prime}-x)/x$, in which $x^{\prime}$ denotes the actual value and $x$ is the ideal value.
@@ III Preparation of high-fidelity entangled cat state
The last term in Eq. (24) is the additional composite drives, where $\delta_{n}=\omega_{n}-\omega_{k}$ is the detuning of the $n$th two-photon drive, and $N$ represents the number of two-photon drives that we may introduce for the stable entangled cat states. According to effective Hamiltonian theory James and Jerke (2007), the effective Hamiltonian in the limit of large detunings can be derived as
@@
where $\frac{1}{\varepsilon_{n}}=({\frac{1}{\Delta}+\frac{1}{\delta_{n}}})$. Similar to the derivation from Eq. (1) to Eq. (10), we can get a new effective Hamiltonian by restricting the Hamiltonian $H_{\mathrm{eff},N}(t)$ to the cat-state subspace as
@@
For simplicity, we define $\frac{4\Omega_{n}}{\varepsilon_{n}}=r_{n}$ and $\zeta=(\Delta-\delta_{n})/n$. Then the effective Hamiltonian $H_{\mathrm{eff},N}^{\prime}(t)$ is expressed as
@@
To understand the construction of composite pulses, we introduce the dimensionless position operator $x=\frac{1}{\sqrt{2}}(a_{0}+a_{0}^{{\dagger}})$ and the momentum operator $p=\frac{i}{\sqrt{2}}(a_{0}^{{\dagger}}-a_{0})$. The Hamiltonian in Eq. (29) is reformulated as
$\mathcal{H}_{\mathrm{MS},N}(t)=f(t)S_{x}x+g(t)S_{x}p,$ (30)
@@
As seen in Eq. (31), the evolution of the operator $\mathcal{U}(t)$ is related to the trajectorys with dimensionless coordinates $\mathcal{F}(t)$ and $\mathcal{G}(t)$ in the $x-p$ phase-space. Equation (32) shows that $\mathcal{F}(t)$ and $\mathcal{G}(t)$ are respectively the integrations of $f(t)$ and $g(t)$ Sørensen and Mølmer (2000); Haddadfarshi and Mintert (2016). Ideally, when the time $t=\tau=\frac{2\pi}{\zeta}$, $\mathcal{F}(\tau)=0$, $\mathcal{G}(\tau)=0$, and the phase-space trajectorys of $(\mathcal{F}(t),\mathcal{G}(t))$ is closed as seen in Fig. 3(a). Meanwhile, the operator $\mathcal{U}(\tau)=e^{-i\mathcal{A}(\tau)S_{x}^{2}}$, that is, only the internal spin evolution is retained. However, errors will lead to the phase-space trajectory not closed completely, which in turn may causes infidelity of entangled cat states Rauschenbeutel et al. (1999); Wang et al. (2021); Webb et al. (2018). As we can see from Fig. 3(b), the circular trajectory becomes more complete with the increase of $N$. In other words, we can reduce errors and improve the fidelity of the scheme by increasing $N$.
@@
In Fig. 4, we plot the populations of the cat states with different $N$ ($N=1,2,4,8$). The curves with $N=1$ correspond to the curves in Fig. 2(a), i.e., the populations of the cat states under the cat-code MS gate Chen et al. (2022). As shown in Fig. 4, compared to the results with $N=1$, the entangled cat states can exist for a longer duration when $N=2,4,8$. This means that more stable entangled cat states can be obtained by simply increasing $N$. Here, the parameters $\{\alpha,K,\Omega_{p},\Delta\}$ are same as in Fig. 2, and $\zeta=\Delta$.
Figure 5: The effective circuit diagram of the protocol. The shaded area which is an array of Josephson junctions represents the KPO, and the $LC$ oscillator represents the cavity mode $\mathcal{A}_{0}$ in our protocol. The KPO coupled to the $LC$ oscillator via capacitor $C_{g}$. Here, $C_{J}$ and $E_{J}$ correspond to the capacitance and energy of the Josephson junctions array, respectively, and $\Phi(t)$ represents the external magnetic flux. $C_{s}$ and $C_{g}$ are two additional large capacitors.
Superconducting quantum interference devices (SQUIDs) can be a possible implementation of the protocol Yang et al. (2003); Yaakobi et al. (2013); Gu et al. (2017); Wang et al. (2019a); You et al. (2007); Flurin et al. (2015); Wustmann and Shumeiko (2013). As shown in Fig. 5, considering the effective circuit of the array of Josephson junctions coupled to an $LC$ oscillator through a large gate capacitance $C_{g}$. The array of Josephson junctions with the energy $E_{J}$ and capacitance $C_{J}$ serves as the KPO, Wallquist et al. (2006); Liao et al. (2010); Arrangoiz-Arriola et al. (2018); Wang et al. (2019b); Xiang et al. (2013), and the $LC$ oscillator serves as the cavity mode $\mathcal{A}_{0}$ with the frequency $\omega_{0}=1/\sqrt{L_{r}C_{r}}$ Wang et al. (2019a); Castellanos-Beltran and Lehnert . The Josephson junction is shunted by an external large capacitance $C_{s}$. Meanwhile, the Josephson energy $E_{J}$ can be adjusted by an external magnetic flux $\Phi(t)$ with frequency $\omega_{p}$, leading to $\tilde{E_{J}}=E_{J}+\delta E_{J}\cos(\omega_{p}t)$.
Utilizing the charge of the circuit and the external magnetic flux, one can construct the initial Hamiltonian that describes the circuit Koch et al. (2007). Then according to the standard quantization procedure of the circuit You et al. (2007); Flurin et al. (2015) and setting appropriate parameters, the Hamiltonian of a single KPO coupled to cavity mode $\mathcal{A}_{0}$ can be derived as Chen et al. (2022); Kang et al. (2022)
$\mathbf{H}=$ $-Ka^{{\dagger}2}a^{2}+\Omega_{p}(a^{{\dagger}2}+a^{2})$ (43)
$+[Jaa_{0}^{\dagger}e^{i\Delta t}+\mathrm{H.c.}].$
The relevant parameters in Eq. (43) can be represented as follows Wang et al. (2019a); Koch et al. (2007); Chen et al. (2022); Kang et al. (2022): $K=2E_{C}/K_{0}^{2}$, $\Omega_{p}=\delta E_{J}\omega_{k}/8E_{J}$, $J=-i2C_{g}eV_{o}n_{0}/(C_{g}+C_{s})$, and detuning $\Delta=\omega_{0}-\omega_{k}$. Where $E_{C}$ is the charging energy of the KPO, $K_{0}$ is the number of SQUIDs, $\omega_{k}=8\sqrt{E_{c}E_{J}/K_{0}}$ is the frequency of KPO, $V_{o}=\sqrt{\frac{\omega_{0}}{2C_{r}}}$ is the root-mean-square voltage of the local oscillator, $n_{0}=\sqrt[4]{E_{J}/(32K_{0}E_{C})}$ is the zero-point fluctuation. Considering the large distance between two KPOs, their coupling interaction can be neglected Wang et al. (2019a); Koch et al. (2007); Chen et al. (2022); Kang et al. (2022). Therefore, the Hamiltonian of two KPOs coupled to a cavity can be derived by summing up Eq. (43).
Figure 6: Fidelities vs parameter errors. (a) timing errors, (b) detuning errors, (c) coupling errors. In figures (a) and (b), increasing $N$ leads to a flatter response to control errors, the robustnesses against parameters errors are enhanced. The other parameters are the same as in Fig. 4.